OpenStudy (anonymous):

Use power series operations to find the Taylor series at x = 0 for cos² x. (Hint: cos² x = (1+ cos2x) / 2.) What is the Taylor series for cos² x at x = 0?

OpenStudy (anonymous):

Recall that \[\cos x=\sum_{k=0}^\infty \frac{(-1)^{k}x^{2k}}{(2k)!}\] This means \[\cos 2x=\sum_{k=0}^\infty \frac{(-1)^{k}(2x)^{2k}}{(2k)!}=\sum_{k=0}^\infty \frac{(-4)^{k}x^{2k}}{(2k)!}\] Add \(1\) and divide by \(2\) and you get the series for \(\cos^2x\).